OpenStudy (anonymous):

Find an equation of the line having the given slope and containing the given point. Express your answer in the form x=a,y=b,or y=mx+b m=-2,(7,0)

OpenStudy (anonymous):

y = -2x + 14

OpenStudy (anonymous):

the slope is -2 and the point they give is (7,0)... y-0=-2(x-7), y=-2x+14